# Towards Mitigating Excessive Forgetting in LLM Unlearning via Entanglement-Aware Unlearning with Proxy Constraint

### v1

@@ Abstract
Large language models (LLMs) are trained on massive datasets that may include private or copyrighted content. Due to growing privacy and ownership concerns, data owners may request the removal of their data from trained models. Machine unlearning provides a practical solution by removing the influence of specific data without full retraining. However, most existing methods lack a sound forgetting boundary, causing some samples to be under-forgotten, leaving residual leakage risks, while others remain over-forgotten at the expense of degraded utility.
In this work, we propose EAGLE-PC (Entanglement-Awareness Guided Loss Reweighting with Proxy Constraint), a novel unlearning framework that addresses these limitations through two key components. First, entanglement-awareness guided loss reweighting determines the forgetting effort of each sample by measuring its similarity to retain samples in the embedding space, enabling more targeted and effective unlearning. Second, a proxy constraint leveraging ICL(In Context Learning)-generated test data softly regularizes the forgetting process, effectively mitigating over-forgetting. EAGLE-PC is compatible with existing gradient-based objectives and serves as a plug-and-play enhancement. We evaluate EAGLE-PC on the TOFU and MUSE benchmarks, showing consistent improvements in the forgetting–utility trade-off across multiple LLMs. Combined with the “NPO+GD” optimizer, it approaches full retraining performance, offering a scalable and robust unlearning solution.
1 Introduction
@@
Large language models (LLMs) [25, 1, 2] demonstrate an extraordinary capacity to absorb and retain knowledge from massive web-scale datasets. However, this memorization capability also brings accompanying privacy risks. Real-world training data often contains private, sensitive, or copyrighted content [20, 14], which may unintentionally resurface at inference time. Such unintended disclosure threatens privacy rights [30], undermines intellectual property protections [27], and even facilitates harmful misuse [17]. Such risks highlight the urgent need to remove specific training data (i.e., forget data) from trained models.
Machine unlearning [5, 4] has emerged as a practical alternative that directly updates model parameters to remove the influence of targeted data, without the need for full retraining. The common paradigm for unlearning is to first formulate a dedicated unlearning objective and then subsequently update model parameters through iterative optimizations [16, 13, 29]. These methods are substantially more efficient than retraining from scratch and enable better trade-offs between effectiveness, efficiency, and performance. Although unlearning reduces the prohibitive cost of retraining, it faces inherent risks of over-forgetting. Overly aggressive parameter updates can cause the model to forget not only the targeted information but also useful knowledge, thereby disrupting the balance between forgetting and utility. Furthermore, state-of-the-art LLMs are carefully aligned with safety guidelines, but excessive updates may inadvertently weaken these safety guardrails, posing additional risks to the model’s alignment. Suck risks can be attributed to two main factors: the heterogeneity of forget samples in terms of memorization and entanglement with the retain dataset, which causes some to be over-forgotten while others are under-forgotten; and the lack of convergence guarantees or explicit stopping criteria in the unlearning process, which can lead to over-forgetting across all forget samples.
@@
Nevertheless, current methods remain inadequate in preventing over-forgetting. Existing strategies that consider only individual sample properties [10, 28] overlook the interactions between forget samples and the retain dataset, which is crucial for precise forgetting and utility preservation. Memorization-score-based method [35] requires repeated retraining, resulting in significant computational overhead and limiting its scalability to large language models. Furthermore, current unlearning objectives provide neither convergence guarantees [16, 33] nor explicit stopping criteria, leaving no principled guidance on when sufficient forgetting has been achieved.
Although the unlearning methods mentioned above have made considerable efforts in balancing forgetting and utility, effectively preventing over-forgetting remains a significant challenge in LLM unlearning. First, measuring the degree of heterogeneous memorization across forget samples is computationally prohibitive, especially for LLMs, and there remains a lack of practical and scalable solutions to model heterogeneous memorization effectively. Secondly, identifying an appropriate stopping criterion remains challenging, as the optimal threshold is highly sensitive to variations in model architectures, dataset distributions, and sample-specific characteristics. This highlights the urgent need for a scalable unlearning method that mitigates over-forgetting for LLM.
@@
In this work, we strive to mitigate over-forgetting in unlearning for LLMs, through entanglement-aware loss reweighting and proxy constraint. We propose EAGLE (Entanglement-Awareness Guided Loss Reweighting), the first method to explicitly model the heterogeneous memorization of forget samples by leveraging their entanglement with the retain dataset to guide adaptive loss reweighting. Specifically, EAGLE measures per-sample similarity in the embedding space and dynamically adjusts the forgetting effort by reweighting the loss for each forget sample accordingly. Given the sensitivity of optimal thresholds to variations in models and data, setting a fixed stopping criteria is unreliable. To better identify an appropriate threshold, we leverage open-source LLMs to construct proxy constraint by predicting retrain model performance on forget samples. For each sample, the LLMS is prompted with a few examples illustrating how the retrained model predicts, and then generates proxy outputs that approximate the retain model’s behavior. These proxy outputs are used in a loss-based penalty mechanism to specify how much the model should still “remember” the targeted information. By penalizing deviations from these proxy targets, the mechanism provides principled guidance to constrain unnecessary loss escalation and prevent over-forgetting.
We conduct comprehensive evaluations on two representative benchmarks: TOFU [18] (using Phi-1.5 and LLaMA2-7B) and MUSE [22] (covering BBC News and Harry Potter subsets with LLaMA2-7B and ICLM). Our EAGLE-PC framework is instantiated on four mainstream unlearning optimizers (GA, GD, NPO, NPO+GD), consistently enhancing unlearning. Notably, EAGLE-PC (NPO+GD) closely matches the performance of the retrain model on TOFU, while EAGLE-PC (GA) even outperforms GD, demonstrating that entanglement-aware guidance using only the average retain embedding can surpass stronger baselines with full retain data. Our contributions are summarized as follows:
  • Entanglement-Awareness Guided Loss Reweighting: We propose EAGLE, a scalable mechanism that reweights the loss contributions of forget samples by quantifying their entanglement with the retain set, measured via embedding similarity, enabling fine-grained and effective unlearning with no additional training.
  • Proxy Constraint: We introduce a soft regularization mechanism based on proxy data synthesized via ICL, which serves as a semantically grounded and adaptive constraint to prevent over-forgetting, effectively balancing forgetting performance and model utility.
@@
  • Unified and Plug-and-Play: EAGLE-PC is general and compatible with existing gradient-based unlearning objectives, including GA, GD, and NPO, enabling plug-and-play improvements without costly retraining.
@@
  • Strong Empirical Evidence: Extensive experiments on two benchmark datasets, TOFU and MUSE, demonstrate that EAGLE-PC achieves finer-grained unlearning with substantially better trade-offs. We validate our approach across multiple models, including Phi, LLaMA2-7B, and ICML-7B. For example, under a 5% unlearning setting with NPO+GD, EAGLE-PC boosts forget quality from 0.0878 to 0.7934. It also preserves high model utility, improving from 0.4158 to 0.4259, and significantly outperforms standard baselines by a wide margin.
We believe our framework provides a practical step toward trustworthy, scalable, and robust machine unlearning for real-world LLM deployment under right-to-be-forgotten demands.
2 Background and Related Work
2.1 Loss Reweighting in Unlearning
Loss reweighting [10, 28] has shown significant benefits for machine unlearning with large language models (LLMs). Huang et.al. [10] proposed a Sample-wise Adaptive Coefficient (inversely proportional to the empirical loss) for Gradient Ascent, which is used to modulate the gradient ascent loss applied to forgetting samples. This approach helps prevent model divergence (or “explosion”) and reduces the update contribution from samples whose influence has already been effectively removed, while prioritizing those with minimal loss, indicating they are not yet sufficiently forgotten. However, as noted in their work, relying solely on empirical loss as an evaluation metric for sample importance is inherently limited and potentially biased. In parallel, [28] identified two distinct objectives for loss reweighting in the context of unlearning: Saturation and Importance. The Saturation criterion emphasizes samples that have not yet been adequately optimized, whereas the Importance criterion focuses on those samples that are most influential for overall loss minimization.
While these two perspectives offer meaningful insights into loss reweighting, their designs focus primarily on the properties of forget samples alone, such as their loss magnitude or optimization status. However, they overlook the interaction with retain samples, which is crucial for balancing utility and forgetting. Ignoring this interaction may lead to over-forgetting or utility degradation. In contrast, we argue that entanglement-awareness guided loss reweighting provides a more principled approach for precise and targeted unlearning.
2.2 Entanglement in Unlearning
Several recent works [3, 35] have investigated the influence of data distribution—particularly the relationship between forget and retain sets—on the effectiveness of unlearning. These studies reveal that the semantic and statistical positioning of forget data relative to the retain knowledge base can significantly affect both forgetting efficiency and model utility. The first line of work [3] systematically explores how the distributional alignment of forget samples affects the unlearning dynamics in large language models (LLMs). The authors demonstrate that unlearning out-of-distribution (OOD) data typically requires more optimization steps, yet achieves a better trade-off between forgetting and generalization. In contrast, unlearning in-distribution (ID) data leads to a rapid degradation in performance, often overshooting the forgetting target. However, no method is proposed to explicitly leverage this insight in a practical unlearning framework. The second study [35] dives deeper into why some forget samples are harder to erase than others. The authors argue that unlearning difficulty is largely governed by two factors: the degree of entanglement and the extent of memorization. To mitigate this, they introduce RUM (Refined-Unlearning Meta-algorithm), which refines the forget set into homogeneous subsets based on memorization scores. While RUM proves effective, it suffers from two key limitations: (1) The computation of memorization scores incurs high computational overhead. (2) Their empirical validation is limited to ResNet-based image classification tasks without exploring the generalization to LLMs.
In contrast, our work is the first to incorporate the entanglement between forget and retain data into the unlearning process, enabling more precise and targeted forgetting while avoiding costly memorization score computations and striking a practical balance between effectiveness and efficiency.
2.3 Over-forgetting in Unlearning
@@
Among existing unlearning methods, Gradient Ascent (GA) is widely recognized for its strong forgetting capability. However, it often leads to over-forgetting, which degrades the model’s performance on unrelated tasks. This limitation is largely attributed to the use of unbounded forget losses and under-representative retain data, which together make it challenging to maintain a proper balance between forgetting and utility preservation. To mitigate this, GA-based variants introduce gradient upper bounds to encourage early convergence. Yet, such adjustments frequently result in under-forgetting, where sensitive or undesirable information is only partially removed [7, 26]. An alternative line of work seeks to reverse the conventional training objective using an assistant LLM. For instance, Ji et al. [12] propose training an auxiliary model with inverted training signals to address both over- and under-forgetting. While promising, this approach comes with significant computational overhead, as training an additional assistant model can be resource-intensive.
3 Preliminary
3.1 Unlearning Problem Formulation
Let $\theta^{o}=\mathcal{A}(\mathcal{D}_{\text{train}})$ be the weights obtained by applying a training algorithm $\mathcal{A}$ on a training dataset $\mathcal{D}_{\text{train}}$. We will refer to $\theta^{o}$ as the “original model”. Further, let $\mathcal{D}_{f}\subseteq\mathcal{D}_{\text{train}}$ denote a subset of the training data referred to as the “forget dataset”. For convenience, we will refer to its complement as the “retain set” $\mathcal{D}_{r}=\mathcal{D}_{\text{train}}\setminus\mathcal{D}_{f}$. Informally, the goal of an unlearning algorithm $\mathcal{U}$ is to utilize $\theta^{o}$, $\mathcal{D}_{f}$ and $\mathcal{D}_{r}$ to produce an unlearned model $\theta^{u}=\mathcal{U}(\theta^{o},\mathcal{D}_{r},\mathcal{D}_{f})$ from which the influence of $\mathcal{D}_{f}$ is removed and retains the other knowledge/capabilities that the original LLM possesses, including $\mathcal{D}_{r}$. A commonly adopted formalism is to compare the distribution of $\theta^{u}$ with that of $\theta^{r}=\mathcal{A}(\mathcal{D}_{\text{train}}\setminus\mathcal{D}_{f})$, i.e., a model retrained from scratch after excluding the forget dataset. The ideal unlearning algorithm, then, according to this viewpoint, is one that yields the same distribution of weights as retraining from scratch.
3.2 Loss Functions for Unlearning

### v2

# Towards Mitigating Excessive Forgetting in LLM Unlearning via Entanglement-Guidance with Proxy Constraint

@@ Abstract.
Large language models (LLMs) are trained on massive datasets that may include private or copyrighted content. Due to growing privacy and ownership concerns, data owners may request the removal of their data from trained models. Machine unlearning provides a practical solution by removing the influence of specific data without full retraining. However, most existing methods still suffer from over-unlearning due to the lack of a principled mechanism to regulate the forgetting boundary, leading to unnecessary utility degradation and heightened privacy and robustness risks.
In this work, we propose EGUP (Entanglement-Guided Unlearning with Proxy Constraint), a novel framework that leverages entanglement and proxy constraint to guide the unlearning process while mitigating over-unlearning. Within each iteration, EGUP employs inter-sample entanglement to adaptively reweight the unlearning strength, assigning greater unlearning efforts to forget samples that are semantically closer to retained knowledge. Across iterations, EGUP leverages intra-sample entanglement to track the representation shift of each forget sample and dynamically adjust its unlearning effort. In addition, we incorporate a proxy constraint that approximates the model’s expected outputs after unlearning, forming a reference boundary that softly regularizes the unlearning process. EGUP is compatible with existing gradient-based objectives and serves as a plug-and-play enhancement. We evaluate EGUP on the TOFU and MUSE benchmarks, demonstrating consistent improvements in the unlearning–utility trade-off across multiple LLMs. Moreover, EGUP achieves performance close to the retrained model while remaining scalable and robust.
* This work is independent of and outside of the work at Amazon
🖂 Corresponding authors.
copyright: none
1. Introduction
@@
Large language models (LLMs) (Touvron et al., 2023; Achiam et al., 2023; Bai et al., 2023) demonstrate an extraordinary capacity to absorb and retain knowledge from massive web-scale datasets. However, this memorization capability also brings accompanying privacy risks. Real-world training data often contains private, sensitive, or copyrighted content (Patil et al., 2023; Karamolegkou et al., 2023), which may not only unintentionally resurface at inference time, but can also be maliciously extracted by adversaries through targeted attacks such as training data extraction and membership inference (Meng et al., 2025; Song et al., 2025). Such disclosures threaten individual privacy rights (Yao et al., 2024b), undermines intellectual property protections (Wei et al., 2024), and may even facilitates harmful misuse (Liu et al., 2023). Such risks highlight the urgent need for mechanisms that can reliably remove specific training data (i.e., forget data) from trained models.
Machine unlearning (Cao and Yang, 2015; Bourtoule et al., 2021) has emerged as a practical alternative that directly updates model parameters to remove the influence of targeted data, without the need for full retraining from scratch. The common paradigm for unlearning is to first formulate a dedicated unlearning objective and then subsequently update model parameters through iterative optimizations (Liu et al., 2025; Jia et al., 2024; Yao et al., 2024a). By explicitly guiding the optimization process toward suppressing memorized information, these approaches significantly reduce computational cost while maintaining model functionality on retained knowledge. Consequently, machine unlearning offers a promising balance between effectiveness, efficiency, and utility, making it particularly suitable for LLM scenarios.
Although unlearning substantially reduces the prohibitive cost of full retraining, it still faces the inherent risk of over-unlearning. When parameter updates become overly aggressive, the model may not only erase the targeted knowledge but also inadvertently degrade useful capabilities, thereby disrupting the desired balance between forgetting and utility (Zhao, 2025). In addition, state-of-the-art LLMs are carefully aligned with safety guidelines; excessive perturbations on model parameters may weaken these safety guardrails, potentially compromising model alignment and increasing the likelihood of generating unsafe content (Łucki et al., 2024; Chen et al., 2025). Beyond utility and safety concerns, recent studies further reveal that over-unlearning can amplify privacy and robustness risks. Excessive representation drift may destabilize decision boundaries and induce abnormal confidence behavior, which in turn increases the success rate of membership inference attacks on both forget and retain samples, while simultaneously reducing robustness under adversarial perturbations (Hayes et al., 2025; Xue et al., 2025). These findings highlight that over-unlearning is not merely a degradation in performance, but a broader threat to model reliability, safety, and security.
@@
Nevertheless, existing unlearning methods remain inadequate in effectively preventing over-unlearning. Approaches that consider only individual sample properties (Huang et al., 2024; Yang et al., 2025) fail to account for the complex interactions between forget samples and the retain dataset, which exert significant influence on the forgetting behavior and are crucial for maintaining a desirable balance between effectiveness and utility preservation. Another line of work leverages memorization scores to guide unlearning (Zhao et al., 2024). However,its reliance on repeated retraining leads to significant computational overhead, making it impractical for LLMs. Furthermore, current unlearning objectives inherently lack reliable termination conditions (Liu et al., 2025; Ye et al., 2025) and explicit stopping criteria, leaving no principled guidance on when sufficient unlearning has been achieved.
Despite these substantial efforts in mitigating over-unlearning, effectively preventing it remains a fundamental and largely unsolved challenge in LLM unlearning. We attribute such risks to two main factors. First, forget samples exhibit strong heterogeneity in memorization and their sematic entanglement with the retain dataset. As a result, applying uniform unlearning effort may cause some samples to be over-unlearned while others remain under-unlearned. Second, most existing work lack reliable convergence guarantees and explicit stopping criteria, making it difficult to determine when adequate unlearning has been achieved, thereby increasing the risk of over-unlearning across all forget samples.
In this work, we propose EGUP (Entanglement Guided Unlearning with Proxy Constraint), an entanglement-guided unlearning framework for large language models that mitigates over-unlearning through entanglement guidance and proxy regularization. EGUP explicitly captures inter- and intra-sample entanglement to guide the unlearning process. Specifically, inter-sample entanglement, defined across samples, captures the semantic coupling between forget and retain samples in the embedding space and leverages this relationship to adaptively modulate unlearning effort across samples via sample-level loss reweighting, enabling more precise and targeted unlearning. In contrast, intra-sample entanglement, defined across iterations, measures the temporal consistency of a forgotten sample’s representations throughout the training process. It also reflects how robust this sample remains to perturbations introduced by updates on other forget samples within the shared parameter space. By serving as a temporal regulation signal, it helps smooth the unlearning trajectory, thereby preventing instability and avoiding overly aggressive unlearning.
@@
Furthermore, the absence of a clear stopping criterion makes it difficult to determine when unlearning is sufficient, increasing the risk of over-unlearning. A naive approach might be to rely on a fixed threshold to terminate unlearning; however, this is unreliable because the optimal stopping point is highly sensitive to variations in model architecture, data distribution, and individual samples. To better address this, we leverage open-source LLMs to construct proxy constraint by predicting retrained model performance on forget samples. For each sample, the LLM is prompted with a few examples illustrating how the retrained model predicts, and then generates proxy outputs that approximate the retain model’s behavior. These proxy outputs are used in a loss-based penalty mechanism to specify how much the model should still “remember” the targeted information. By penalizing deviations from these proxy targets, the mechanism provides principled guidance to constrain unnecessary loss escalation and prevent over-unlearning.
We extensively evaluate EGUP on two representative benchmarks, TOFU (Maini et al., 2024) and MUSE (Shi et al., 2024). When integrated into diverse mainstream unlearning pipelines, EGUP consistently outperforms baselines and further achieves almost the same performance as full retraining. More importantly, EGUP consistently outperforms baselines that depend on retain data, while using only the averaged retain embedding. Our contributions are summarized as follows:
  • Entanglement-Guided Unlearning Mechanism: We design an entanglement-guided unlearning framework that jointly leverages inter-sample and intra-sample entanglement to regulate the unlearning process. By reweighting sample-level unlearning effort based on semantic entanglement and incorporating a temporal regularization term sensitive to representation drift, the framework achieves more precise and stable unlearning.
  • Proxy Constraint: We introduce a soft regularization mechanism based on proxy outputs synthesized via in-context learning (ICL) to approximate a retain model’s behavior on forget samples, providing a principled constraint that prevents over-unlearning.
@@
  • Unified and Plug-and-Play: EGUP is general and compatible with multiple existing gradient-based unlearning objectives, enabling plug-and-play improvements without costly retraining.
@@
  • Strong Empirical Evidence: Extensive experiments on TOFU and MUSE show that EGUP achieves superior trade-offs across models, including Phi-1.5, LLaMA2-7B, and ICML-7B. Under a 5% unlearning setting with NPO+GD, it boosts forget quality from 0.0878 to 0.7934 while maintaining utility, substantially outperforming standard baselines.
2. Background and Related Work
2.1. Unlearning for LLMs
Machine unlearning aims to remove the influence of specific training data from a trained model while preserving overall performance on the retained data. Recent works have introduced gradient-based optimization strategies to achieve this goal. Notably, Gradient Difference (GD) (Maini et al., 2024) and Negative Preference Optimization (NPO) (Zhang et al., 2024) represent two seminal approaches that guide parameter updates to counteract the influence of forget samples while maintaining model stability. Building upon these methods, subsequent studies refine the underlying loss functions, introducing adaptive weighting and optimization mechanisms to achieve more stable and precise unlearning. Huang et al. (Huang et al., 2024) proposed a Sample-wise Adaptive Coefficient, inversely scaled by empirical loss, to balance gradient ascent on forget samples. It prevents redundant updates to already forgotten data while emphasizing those that remain insufficiently erased. However, as noted in their work, relying solely on empirical loss as a measure of sample importance introduces bias and fails to capture representational complexity. In parallel, Yang et al. (Yang et al., 2025) identify two complementary criteria for loss reweighting in unlearning: Saturation, which prioritizes under-optimized samples, and Importance, which emphasizes samples most influential for overall loss minimization. Together, these methods demonstrate the potential of adaptive weighting schemes to balance unlearning quality and model retention.
2.2. Memorization
Deep neural networks are known to memorize their training data. Feldman et al. (Feldman, 2020a) show that such instance-level memorization is not merely an artifact of overfitting but a key mechanism for generalization under long-tailed distributions. Subsequent studies (Feldman, 2020a; Brown et al., 2021; Attias et al., 2024) further show that moderate memorization is necessary to capture rare or low-frequency patterns critical for generalization.
Beyond its role in generalization, recent research (Torkzadehmahani et al., 2024) reveals that memorization and unlearning are deeply intertwined. At the extreme, non-memorized examples can be regarded as trivially unlearned, since their predictions would remain unchanged without being included in training. In contrast, highly memorized examples are much harder to erase, as their influence is more deeply embedded in the model parameters. Zhao et al. (Zhao et al., 2024) systematically analyze the role of memorization in unlearning difficulty. They find that approximate unlearning methods are generally more successful on forget sets with low-memorization samples, while highly memorized ones are substantially harder to remove. To address this, they propose RUM (Refined-Unlearning Meta-algorithm), which partitions the forget set into homogeneous subsets based on memorization scores, achieving more stable and fine-grained unlearning. Despite its strengths, RUM still faces several limitations: (1) The computation of memorization scores incurs high computational overhead. (2) It focuses only on inter-sample correlations, ignoring intra-sample representational dynamics essential for modeling temporal entanglement and stable unlearning. (3) Their empirical validation is limited to ResNet-based image classification tasks without exploring the generalization to LLMs.
2.3. over-unlearning in Unlearning
Continued gradient updates during unlearning can excessively alter the model’s behavior on forget samples, causing their predictions to diverge from those on naturally unseen inputs. Such over-unlearning not only degrades generalization and overall model utility but can also introduce new privacy risks. Such over-unlearning not only degrades generalization and model utility but also increases vulnerability to Membership Inference Attacks (MIA) (Shokri et al., 2017), potentially leaking whether a sample in the forget dataset $\mathcal{D}_{f}$ was part of the original training set $\mathcal{D}$ (Shi et al., 2024; Hayes et al., 2025; Pawelczyk et al., 2024).
@@ 2.2. Memorization
To address these challenges, Huang et al. (Huang et al., 2024) incorporate gradient upper bounds to stabilize the optimization process and facilitate early convergence. However, it can inadvertently result in under-unlearning of sensitive information. An alternative line of work seeks to reverse the conventional training objective using an assistant LLM. For instance, Ji et al. (Ji et al., 2024) propose training an auxiliary model with inverted training signals to address both over- and under-unlearning. While promising, this approach comes with significant computational overhead, as training an additional assistant model can be resource-intensive.
3. Preliminary
3.1. Unlearning Problem Formulation
In machine unlearning, let $\mathcal{D}=\{z_{i}\}_{i=1}^{N}$ represent a pretraining dataset of $N$ data points, including $z_{i}=(x_{i},y_{i})$ features and labels in supervised learning. The training dataset $\mathcal{D}$ comprises two subsets, namely $\mathcal{D}_{f}$ and $\mathcal{D}_{r}$, representing the data points to be unlearned and the retain data points, respectively, wherein $\mathcal{D}_{r}=\mathcal{D}\setminus\mathcal{D}_{f}$. Denote $\mathcal{M}_{t}=\mathcal{M}(\mathcal{D})$ as the model trained on the entire dataset $\mathcal{D}$. The goal of MU is to produce an unlearned model $\mathcal{M}_{u}$ that closely resembles the model $\mathcal{M}_{u}^{*}$ trained solely on $\mathcal{D}_{r}$, i.e., $\mathcal{M}_{u}^{*}=\mathcal{M}(\mathcal{D}_{r})$.
3.2. Mathematical Modeling for Unlearning
A mainstream approach in current LLM unlearning research is to fine-tune the pre-trained model using specially designed loss functions that promote unlearning. While the specific formulations differ across methods, these objectives typically share an underlying structure as follows:
(1) $\min_{\theta^{\prime}}L(\theta^{\prime})=\min-L(\theta^{\prime},\mathcal{D}_{f})+\alpha L(\theta^{\prime},\mathcal{D}_{r}),$
where $\alpha$ is a hyper-parameter for controlling the retain strength. The first loss term, $L(\theta^{\prime},\mathcal{D}_{f})$, which we call the forget loss, measures the prediction quality on the forget dataset. A typical choice of the forget loss is the (next-token prediction) cross-entropy loss on the forget dataset. The second loss term $L(\theta^{\prime},\mathcal{D}_{r})$, which we call the retain loss, measures the prediction quality on the retain dataset $\mathcal{D}_{r}$. Equation 1 essentially maximizes the forget loss while minimizing the retain loss, so this objective should ideally simultaneously achieve the aforementioned two goals.